You have an infrared laser that produces light of an unknown wavelength.
A. You shine this laser through a mask containing a single slit of width a. On a screen that is a
distance L away, you find a very broad bright spot at the center of the screen that has a width of
L. Determine the wavelength of the laser in terms of the given quantities.
B. You shine the same laser through a different mask containing a single slit that
1
Expert's answer
2020-12-04T12:18:26-0500

As per the question,

Slit width = a

Distance between the screen and slit = L

We know that width of the central maxima = "\frac{2\lambda D}{d}"

Now substituting the values as per the given question,

"L=\frac{2\lambda L}{a}"

Hence, "\lambda = \frac{a}{2}"

and we know that "x=\frac{2\lambda D}{d}"

Hence, "x\propto \frac{1}{d}"